A fuel oil heater receives 23200 lb of oil per hour at 80°F and discharges it at 160°F. The specific heat of the oil is 0.48 BTU/lb-°F. Find the rate of heat transfer to the oil, BTU/hr.
1
Expert's answer
2020-03-13T11:14:40-0400

As per the given question,

quantity of flue=23200 lb

"T_1=80^oF"

"T_2=160\u00b0F."

specific heat of the oil is "(s)=" 0.48 BTU/lb-°F

we know that heat, "Q=ms(T_2-T_1)"

"Q=232200\times 0.48\times (160-80)=8916480BTU"

Hence rate of heat flow per hour "\dfrac{dQ}{dt}=8916480BTU/hour"